# Estimating Treatment Effect Heterogeneity in Psychiatry: A Review and Tutorial with Causal Forests

### v1

@@ 4 Application: Estimating Resilience
We illustrate the approach by focusing on a recent application of grf to estimate differential resilience to the effects of combat stress in leading to post-traumatic stress disorder (PTSD) among US Army soldiers (Kessler et al., under review). PTSD is the signature mental disorder of war (Paulson and Krippner,, 2007). The roughly 7% of the US population made up of military personnel and veterans are estimated to account for nearly 20% of all cases of PTSD in the US, with an annual societal cost estimated at more than \$230B (DeAngelis,, 2023). However great variation is thought to exist in the effects of combat stress on PTSD among service members (Karstoft et al.,, 2015; Schultebraucks et al.,, 2021). We illustrate the value of causal forests by investigating these effects in a secondary analysis of a sample of US Army soldiers in three combat brigades who participated in a self-reported survey shortly before and then again after returning from a combat deployment in Afghanistan in 2012-2014 as part of the Army Study to Assess Risk and Resilience in Servicemembers (Army STARRS, Ursano et al., (2014)). Exposure to combat stress was assessed in the follow-up surveys administered to this sample after they returned from deployment. As expected, high combat stress exposure was found to be associated significantly with elevated prevalence of current PTSD as assessed in the follow-up surveys (Kessler et al., under review).
The effects of exposure to high combat stress can be approximated by defining a dichotomous measure of the extent of combat stress exposure in the sample who had combat arms occupations (e.g., infantry, artillery). The extent to which information about individual differences in both baseline survey reports (including reports about history of and recent symptoms of PTSD) and baseline administrative data predicted this variation in exposure to high combat stress can then be investigated. The net association of high combat stress with the subsequent occurrence of meeting diagnostic criteria for PTSD in the follow-up survey can then be examined to estimate ATE as well as to inspect the distribution of estimated CATE. More details on the dataset (Papini et al.,, 2023) and a more extensive substantive analysis of the data (Kessler et al., under review) are presented elsewhere.
@@
and so our average treatment effect measures the fraction of healthy soldiers when no one is deployed minus the counterfactual fraction of healthy soldiers when all soldiers are deployed, which we expect to be larger than zero,
$\begin{split}\tau&=\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\right]\\ &=\mathbb{P}\left[Y_{i}(\text{Low combat stress})=\text{Healthy}\right]-\mathbb{P}\left[Y_{i}(\text{High combat stress})=\text{Healthy}\right].\\ &=\mathbb{P}\left[\text{Low resilience}\right].\end{split}$
@@
This type of taxonomy gives us two approaches to measuring resilience: a causal measure based on CATEs (low vs. high resilience); and a predictive measure based on risk (high vs. low risk). How these two approaches differ in practice is an empirical question we address in Section 4.6.
@@ 4.1 An Exploratory Analysis
We consider 2,542 US soldiers with combat service where 1,542 experienced low combat stress and 1,000 soldiers experienced high combat stress (for details on the sample construction, including details on treatment and outcome definitions we refer to Kessler et al., under review). In the high combat stress group, 96 soldiers developed PTSD post-deployment, while in the low combat stress group, 36 soldiers developed PTSD. The simple difference-in-means (i.e., the mean number of healthy soldiers in the low-stress group minus the mean number of healthy soldiers in the high-stress group) is 7.4% (standard error of 0.9%), and yields a measure of the raw association between deployment and PTSD. This is, however, not necessarily a valid estimate of the ATE, as combat stress exposure during deployment might not be completely random.
@@
A necessary condition for this to be a viable strategy for estimating causal effects is that the propensity score as a function of these variables is bounded away from zero and one, i.e. that we have treatment overlap. Figure 1 shows a histogram of the forest-estimated propensity scores and indicates the overlap assumption is plausible as most of the propensity scores fall within the range 0.5-0.7, that is, along our observable adjustment variables the different soldiers are more or less equally likely to be in either treatment group.
@@ 4.2 Treatment Effect Heterogeneity
The average-case analysis above established that there are a substantial number of low-resilience soldiers in the sample. However, there might be between-individual variation in the effects of high combat stress such that prevalence of PTSD could be reduced by taking this individual variation into account. In order to take a data-driven approach to discover potential subgroups of soldiers that respond differently to combat stress, we divide the dataset into two random groups. In the training set (60% of the data) we fit a CATE estimator, and on a held-out test set (40% of the data) we evaluate these CATE estimates (the exact train/test ratio is more or less a heuristic, we prefer to have slightly more data for training than evaluation in this small dataset).
@@
Figure 3 shows the estimated TOC curve for the training set predicted CATEs, and indicates that there are signs of treatment effect heterogeneity. For example, the top $q=$ 20% of soldiers with the largest predicted treatment effects, have an ATE that is around 11% higher than the overall ATE (with a standard error of $\approx$ 5%). That is, the group of soldiers belonging to the top quintile of predicted treatment effects, had low resilience at an 11% higher rate than average. The confidence intervals are wider at lower quantiles than at higher quantiles since the number of observations is smaller. In this setting, the lowest quantiles of the TOC curve cover zero, highlighting there is considerable estimation uncertainty for this smaller group. We can translate the visual stratification exercise underlying the TOC curve into a single-point estimate by forming an estimate of the area under this curve.
@@
Besides serving as a useful test for heterogeneity, the AUTOC can also be used for model selection. For example, in choosing among candidate predictor models (for example two or more alternative CATE models), the one with the largest AUTOC is the one that most successfully manages to stratify the test set sample. As mentioned in Section 3, causal forests are a two-step algorithm where the first step accounts for confounding with respect to some specified baseline predictor variables via an estimated propensity score as well as baseline effects via conditional mean estimates, and the second step fits a CATE function based on some set of baseline predictor variables. In this application, compared to our sample size, the set of predictor variables considered was large (410), so we began the analysis by using a simple heuristic to reduce the number of variables to help our forest capture relevant heterogeneity: For the CATE estimate, we only used predictor variables that, in the conditional mean model, had variable importance in the top 25-th percentile. grf uses a simple assessment of variable importance that counts how often a particular variable is used for recursive partitioning throughout the forest. The AUTOC for this restricted CATE model was very similar to that for the full predictor set, indicating that this subset of predictors captured parsimoniously the meaningful heterogeneity associated with the full predictor set; this approach is closely related to evaluating binary classifiers with the ROC curve in a predictive analysis. Section 4.6 gives one more example of using the AUTOC for evaluating treatment targeting strategies.
Figure 3: Targeting Operator Characteristic curve for causal forest CATE predictions, estimated on a held-out test set. The dashed lines are pointwise 95% confidence intervals that are conditional on our estimated CATE function, and quantify test set uncertainty in estimating the TOC.
@@ 4.3 Evaluating treatment rules with Qini curves
Figure 4 shows the Qini curve for this case study based on our estimated CATE function for our sample of $n=$ 2,542 soldiers. The straight dashed line in Figure 4 is a non-targeting baseline policy that shows the reduction in total PTSD cases we could expect by intervening randomly with various numbers of soldiers. The far right of this curve equals $n\cdot\widehat{\text{ATE}}\approx 180$ soldiers, the PTSD cases avoided if no one is deployed. The solid black curve shows results that can be achieved by using our CATE estimate to target intervention. For example, if we use causal forests to choose 500 soldiers predicted to benefit most, we estimate an overall reduction in PTSD cases of around 90 cases (standard error of 25). To achieve the same benefit by intervening randomly we would have to hold back 1,300 soldiers.
@@ 4.4 Describing estimated heterogeneity via covariate profiles
Figure 5: Histograms of 6 predictor variables for soldiers that are in the bottom and top quartile of predicted test set CATEs (Figure 2(b)). These 6 predictor variables have high variable importance as measured by the causal forest split frequency. The variables are: “z_anx_30d_sx_T0” (anxiety symptoms), “z_dep_30d_sx_T0” (depression symptoms), “z_negativeaffect” (negative affectivity), “z_neurophys_30d_T0” (neuro/physiological symptoms), “z_ptsd_30d_sx_T0” (PTSD symptoms), and “z_sum_30d_sx_T0” (all symptoms). All variables are constructed as a standardized continuous score for 30-day symptoms.
@@ 4.5 Inference on effect modifiers via best linear projections
Formally, we have $\{\alpha^{*},\,\beta^{*}\}=\operatorname{argmin}_{\alpha,\beta}\left\{\mathbb{E}\left[(\tau(X_{i})-\alpha-Z_{i}\beta)^{2}\right]\right\}$. This is not a standard linear regression problem as $\tau(X_{i})$ is not observable. However, grf enables inference about $\beta$ via a built-in AIPW-style construction that has similar statistical properties as what one would expect of classical parametric summaries and can be used to justify exact confidence intervals (Semenova and Chernozhukov,, 2021). In forming estimates like these typical considerations that arise in the familiar parametric setting apply concerning multiple testing and pre-specified hypotheses. Note that, when we deploy a best linear projection we are not assuming that the data is linear, we are simply summarising a non-linear object, the CATE, via a linear approximation.
@@ 4.6 Risk-based targeting and treatment rule model selection
Figure 6(a) shows a histogram of this estimated probability for soldiers in the held-out test sample (where we used all 410 pre-treatment characteristics as predictor variables). The question is, do these risk estimates perform well in targeting soldiers with different treatment benefits? Our intuition is that soldiers that have a large predicted risk of developing PTSD, also have a large treatment benefit of intervention. Using the same quartile stratification exercise as in Figure 2(b), we could stratify the sample by risk quartiles, and then estimate average treatment effects. However, while visually appealing, our end goal is to compare targeting strategies and comparing quartile figures is not as convenient for this purpose. To this end, it turns out we can use the TOC construction to facilitate a side-by-side comparison.
@@
where the key point to recognize is that it is the relative ranking of the estimated CATEs that make up this curve. This means we can use this same construction to evaluate other intervention strategies as long as they imply a ranking/prioritization. Using our estimated risk model to prioritize soldiers with high probabilities of PTSD, we get the TOC:
@@ 5 Discussion
We have surveyed how causal forests can be used to generate both principled estimates of ATE in observational studies and to detect treatment heterogeneity in either experimental or observational studies. For more details on grf, including additional functionality such as missing values support, tree-based policy learning, loss to-follow-up corrections, and more, we refer to vignettes and journal references on the grf website at grf-labs.github.io/grf. As noted in the introduction, there are also a number of alternative machine learning approaches for CATE estimation that are popular in practice (e.g., Athey and Imbens,, 2016; Hahn et al.,, 2020; Kennedy,, 2023; Künzel et al.,, 2019; Luedtke and van der Laan,, 2016; Nie and Wager,, 2021; Tian et al.,, 2014). Customizable Python libraries that implement several of these methods include causalML (Chen et al.,, 2020) and econML (Battocchi et al.,, 2019); and the software toolkit for machine-learning based causal inference is still rapidly growing.

### v2

@@ 4 Application: Estimating Resilience
We illustrate the approach by focusing on a recent application of grf to estimate differential resilience to the effects of combat stress in leading to post-traumatic stress disorder (PTSD) among US Army soldiers (Kessler et al.,, 2024). PTSD is the signature mental disorder of war (Paulson and Krippner,, 2007). The roughly 7% of the US population made up of military personnel and veterans are estimated to account for nearly 20% of all cases of PTSD in the US, with an annual societal cost estimated at more than \$230B (DeAngelis,, 2023). However great variation is thought to exist in the effects of combat stress on PTSD among service members (Karstoft et al.,, 2015; Schultebraucks et al.,, 2021). We illustrate the value of causal forests by investigating these effects in a secondary analysis of a sample of US Army soldiers in three combat brigades who participated in a self-reported survey shortly before and then again after returning from a combat deployment in Afghanistan in 2012-2014 as part of the Army Study to Assess Risk and Resilience in Servicemembers (Army STARRS, Ursano et al., (2014)). Exposure to combat stress was assessed in the follow-up surveys administered to this sample after they returned from deployment. As expected, high combat stress exposure was found to be associated significantly with elevated prevalence of current PTSD as assessed in the follow-up surveys (Kessler et al.,, 2024).
The effects of exposure to high combat stress can be approximated by defining a dichotomous measure of the extent of combat stress exposure in the sample who had combat arms occupations (e.g., infantry, artillery). The extent to which information about individual differences in both baseline survey reports (including reports about history of and recent symptoms of PTSD) and baseline administrative data predicted this variation in exposure to high combat stress can then be investigated. The net association of high combat stress with the subsequent occurrence of meeting diagnostic criteria for PTSD in the follow-up survey can then be examined to estimate ATE as well as to inspect the distribution of estimated CATE. More details on the dataset (Papini et al.,, 2023) and a more extensive substantive analysis of the data (Kessler et al.,, 2024) are presented elsewhere.
@@
and so our average treatment effect measures the fraction of healthy soldiers when no one is deployed minus the counterfactual fraction of healthy soldiers when all soldiers are deployed, which we expect to be larger than zero.
$\begin{split}\tau&=\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\right]\\ &=\mathbb{P}\left[Y_{i}(\text{Low combat stress})=\text{Healthy}\right]-\mathbb{P}\left[Y_{i}(\text{High combat stress})=\text{Healthy}\right]\\ &=\mathbb{P}\left[\text{Low resilience}\right].\end{split}$
@@
This type of taxonomy gives us two approaches to measuring resilience: a causal measure based on CATEs (low vs. high resilience) and a predictive measure based on risk (high vs. low risk). How these two approaches differ in practice is an empirical question we address in Section 4.6.
@@ 4.1 An Exploratory Analysis
We consider 2,542 US soldiers with combat service, where 1,542 experienced low combat stress, and 1,000 soldiers experienced high combat stress (for details on the sample construction, including details on treatment and outcome definitions, we refer to Kessler et al., (2024)). In the high combat stress group, 96 soldiers developed PTSD post-deployment, while in the low combat stress group, 36 soldiers developed PTSD. The simple difference-in-means (i.e., the mean number of healthy soldiers in the low-stress group minus the mean number of healthy soldiers in the high-stress group) is 7.4% (standard error of 0.9%), and yields a measure of the raw association between deployment and PTSD. This is, however, not necessarily a valid estimate of the ATE, as combat stress exposure during deployment might not be completely random.
@@
A necessary condition for this to be a viable strategy for estimating causal effects is that the propensity score as a function of these variables is bounded away from zero and one, i.e., we have treatment overlap. Figure 2 shows a histogram of the forest-estimated propensity scores and indicates the overlap assumption is plausible as most of the propensity scores fall within the range 0.5-0.7, that is, along our observable adjustment variables the different soldiers are more or less equally likely to be in either treatment group.
@@ 4.2 Treatment Effect Heterogeneity
The average-case analysis above established that there are a substantial number of low-resilience soldiers in the sample. However, there might be between-individual variation in the effects of high combat stress such that prevalence of PTSD could be reduced by taking this individual variation into account. In order to take a data-driven approach to discover potential subgroups of soldiers that respond differently to combat stress, we divide the dataset into two random groups. In the training set (60% of the data), we fit a CATE estimator, and on a held-out test set (40% of the data), we evaluate these CATE estimates (the exact train/test ratio is more or less a heuristic; we prefer to have slightly more data for training than evaluation in this small dataset).
@@
Figure 4: Targeting Operator Characteristic curve for causal forest CATE predictions, estimated on a held-out test set. The dashed lines are pointwise 95% confidence intervals that are conditional on our estimated CATE function and quantify test set uncertainty in estimating the TOC.
Figure 4 shows the estimated TOC curve for the training set predicted CATEs, and indicates that there are signs of treatment effect heterogeneity. For example, the top $q=$ 20% of soldiers with the largest predicted treatment effects have an ATE that is around 11% higher than the overall ATE (with a standard error of $\approx$ 5%). That is, the group of soldiers belonging to the top quintile of predicted treatment effects had low resilience at an 11% higher rate than average. The confidence intervals are wider at lower quantiles than at higher quantiles since the number of observations is smaller. In this setting, the lowest quantiles of the TOC curve cover zero, highlighting there is considerable estimation uncertainty for this smaller group. We can translate the visual stratification exercise underlying the TOC curve into a single-point estimate by forming an estimate of the area under this curve.
@@
Beyond serving as a valuable test for heterogeneity, the AUTOC can also assist in model selection. For instance, when comparing candidate predictor models (e.g., two or more alternative CATE models), the model with the highest AUTOC is the one that most effectively stratifies the test set sample. As mentioned in Section 3, causal forests are a two-step algorithm where the first step accounts for confounding with respect to some specified baseline predictor variables via an estimated propensity score as well as baseline effects via conditional mean estimates and the second step fits a CATE function based on some set of baseline predictor variables. In this application, compared to our sample size, the set of predictor variables considered was large (410), so we began the analysis by using a simple pre-screening heuristic to reduce the number of variables to help our forest capture relevant heterogeneity (Athey and Wager,, 2019): For the CATE model, we only used predictor variables that, in the conditional mean model, had variable importance in the top 25-th percentile. The AUTOC for this restricted CATE model was very similar to that for the full predictor set, indicating that this subset of predictors captured parsimoniously the meaningful heterogeneity associated with the full predictor set; this approach is closely related to evaluating binary classifiers with the ROC curve in a predictive analysis. Section 4.6 gives one more example of using the AUTOC for evaluating treatment targeting strategies.
Remark 2.
grf uses a simple assessment of variable importance, first discussed by Breiman, (2001), that counts how often a particular variable is used for recursive partitioning throughout the forest. For a broader discussion of alternative variable importance measures with forests see Bénard and Josse, (2023) and Hastie et al., (2009).
@@ 4.3 Evaluating treatment rules with Qini curves
Figure 5 shows the Qini curve for this case study based on our estimated CATE function for our sample of $n=$ 2,542 soldiers. The straight dashed line in Figure 5 is a non-targeting baseline policy that shows the reduction in total PTSD cases we could expect by intervening randomly with various numbers of soldiers. The far right of this curve equals $n\cdot\widehat{\text{ATE}}\approx 180$ soldiers, the PTSD cases avoided if no one is deployed. The solid black curve shows results that can be achieved by using our CATE estimate to target intervention. For example, if we use causal forests to choose 500 soldiers predicted to benefit most, we estimate an overall reduction in PTSD cases of around 90 cases (standard error of 25). To achieve the same benefit by intervening randomly, we would have to hold back 1,300 soldiers.
@@ 4.4 Describing estimated heterogeneity via covariate profiles
Figure 6: Histograms of 6 predictor variables for soldiers that are in the bottom (resilient) and top quartile (less resilient) of predicted test set CATEs (Figure 3(b)). These 6 predictor variables have high variable importance as measured by the causal forest split frequency. The variables are: “z_anx_30d_sx_T0” (anxiety symptoms), “z_dep_30d_sx_T0” (depression symptoms), “z_negativeaffect” (negative affectivity), “z_neurophys_30d_T0” (neuro/physiological symptoms), “z_ptsd_30d_sx_T0” (PTSD symptoms), and “z_sum_30d_sx_T0” (all symptoms). All variables are constructed as a standardized continuous score for 30-day symptoms.
@@ 4.5 Inference on effect modifiers via best linear projections
Formally, we have $\{\alpha^{*},\,\beta^{*}\}=\operatorname{argmin}_{\alpha,\beta}\left\{\mathbb{E}\left[(\tau(X_{i})-\alpha-Z_{i}\beta)^{2}\right]\right\}$. This is not a standard linear regression problem as $\tau(X_{i})$ is not observable. However, grf enables inference about $\beta$ via a built-in AIPW-style construction that has similar statistical properties as what one would expect of classical parametric summaries and can be used to justify exact confidence intervals (Semenova and Chernozhukov,, 2021). In forming estimates like these, typical considerations that arise in the familiar parametric setting apply concerning multiple testing and pre-specified hypotheses. Note that, when we deploy a best linear projection we are not assuming that the data is linear, we are simply summarising a non-linear object, the CATE, via a linear approximation.
@@ 4.6 Risk-based targeting and treatment rule model selection
Figure 7(a) shows a histogram of this estimated probability for soldiers in the held-out test sample (where we used all 410 pre-treatment characteristics as predictor variables). The question is, do these risk estimates perform well in targeting soldiers with different treatment benefits? Our intuition is that soldiers that have a large predicted risk of developing PTSD, also have a large treatment benefit of intervention. Using the same quartile stratification exercise as in Figure 3(b), we could stratify the sample by risk quartiles and then estimate average treatment effects. However, while visually appealing, our end goal is to compare targeting strategies, and comparing quartile figures is not as convenient for this purpose. To this end, it turns out we can use the TOC construction to facilitate a side-by-side comparison.
@@
where key point to recognize is that the relative ranking of the estimated CATEs makes up this curve. This means we can use this same construction to evaluate other intervention strategies as long as they imply a ranking/prioritization. Using our estimated risk model to prioritize soldiers with high probabilities of PTSD, we get the TOC:
@@
Remark 3.
We estimated a risk model directly using a random forest with all available predictor variables. In some cases, it may be possible to get better performance by first pre-screening pre-treatment characteristics using general variable selection tools, such as the Lasso. For a practical guide to developing risk models in clinical applications, see Steyerberg, (2019).
@@ 5 Discussion
We have surveyed how causal forests can be used to generate both principled estimates of ATE in observational studies and to detect treatment heterogeneity in experimental or observational studies. For more details on grf, including additional functionality such as missing values support, tree-based policy learning, loss-to-follow-up corrections, and more, we refer to vignettes and journal references on the grf website at grf-labs.github.io/grf.
We have also outlined a general framework for analyzing and comparing output from causal machine learning algorithms on a level playing field with tools such as TOC and Qini curves. As noted in the introduction, there are a number of alternative machine learning approaches for CATE estimation that are popular in practice (e.g., Athey and Imbens,, 2016; Hahn et al.,, 2020; Kennedy,, 2023; Künzel et al.,, 2019; Luedtke and van der Laan,, 2016; Nie and Wager,, 2021; Tian et al.,, 2014). Customizable Python libraries that implement several of these methods include causalML (Chen et al.,, 2020) and econML (Battocchi et al.,, 2019); and the software toolkit for machine learning-based causal inference is still rapidly growing. In practice, it can sometimes be a good idea to try multiple different machine learning-based approaches to CATE estimations; the TOC, Qini curves, and related tools can then be used to compare and benchmark their performance.
@@ Acknowledgments
Conflict of Interest Disclosures: The study protocol was approved by the Research Ethics Committees of the Henry Jackson Foundation and Harvard Medical School (IRB15-0765) with a waiver of informed consent based on data being de-identified. Research has been performed in accordance with the Declaration of Helsinki.